The sum of two numbers is 13, and the sum of their squares is 89. If x is the

first number. Find this number.


1
Expert's answer
2020-11-24T11:36:08-0500

if x is the first number then (13-x) is the second one

So, sum of their squares is 89; x2+(13-x)2=89; x2+169- 26x+x2=89; 2x2-26x+80=0;

x2 -13x+40=0;

D=169-160=9;

x₁=(13+3)/2=8;

x₂=(13-3)/2=5.

checkup: 8+5=13; 64+25=89

So x can be either 5 or 8 .